A force of 10.5N acts on a particle along a direction making an angle 37o with the vertical. Find the component in the vertical direction.

Grade:11

1 Answers

The Hounds Of Justice
5 years ago
The component of the force in the vertical direction will be :
Fperpendicular = F cos\Theta= 10.5 * cos37o
                                  =10.5 * (4/5) = 8.40 N